z^2+9=10z
and
n^2+n=72

Answer by stanbon(57299)


z^2+9=10z
z^2-10z+9=0
(z-9)(z-1)=0
z=9 or z=1
--------------------
and
n^2+n=72
n^2+n-72=0
(n+9)(n-8)=0
n=-9 or n=8
===========
Cheers,
Stan H.